# Incentivizing supplemental math assignments and using AI-generated hints improve exam performance, especially for racially minoritized students

### v1

@@ Abstract
Inequities in student access to trigonometry and calculus are often associated with racial and socioeconomic privilege, and are often propagated to introductory physics course performance. To mitigate these disparities in student preparedness, we developed a two-pronged intervention consisting of (1) incentivized supplemental mathematics assignments and (2) AI-generated learning support tools in the forms of optional hints embedded in the physics assignments. Both interventions are grounded in the Situated Expectancy-Value Theory of Achievement Motivation, which posits that students are more likely to complete a task that they expect to do well in and whose outcomes they think are valuable. For the supplemental math assignments, the extra credit available was scaled to make it worth more points for the students with lower exam scores, thereby creating even greater value for the students who might most benefit from the assignments. For the AI-generated hints, these were integrated into the homework assignments, thereby reducing or eliminating the cost to the student, in terms of time, energy, and social barriers or fear of judgment. Our findings indicate that both these interventions are associated with increased exam scores; in particular, the scaled extra credit reduced disparities in completion of math supplemental material and was associated with reducing racial disparities in exam scores. These interventions, which are relatively simple for any instructor to implement, are therefore very promising for creating more equitable undergraduate quantitative-based courses.
@@ I Introduction
Students’ mathematical skills play an important role in shaping their learning and performance in undergraduate Introductory Physics courses [1, 2, 3, 4]. However, students enter these courses with wide variation in their prior exposure to advanced mathematics (i.e. trigonometry and calculus). Moreover, student access to advanced math courses is associated with racial and socioeconomic privilege, so that Black, Hispanic/Latinx, and Native American students, and students from low-income backgrounds are less likely to have had the opportunity to take advanced math courses in high school [5, 6, 7, 8].
@@
One strategy to address the inequity in prior access to advanced mathematics courses is to provide supplemental instruction to students. This has been found to help students with less math preparation successfully complete Introductory Physics courses [15, 16, 4]. Most supplemental instruction takes the form of an optional in-person course or series of workshops focused on students’ problem solving skills in small groups [17, 16, 15]. This can be quite time and resource intensive. Using online learning tools offers an alternative less resource-intensive way to approach supplemental instruction but relies more heavily on student motivation and might require students to pay additional fees [4, 18]. For example, Forrest et al. [4] offered an optional asynchronous online math tutorial to Introductory Physics students who scored less than 65% on a Math diagnostic exam and found that students who completed the online tutorial were four times more likely to pass the course. However, this tutorial was not free and only about half of the students completed the online tutorial, even though completing the tutorial was associated with a small amount of course points [4]. Thus, motivating students to participate in optional online supplemental math instruction can be challenging [19, 18].
@@
An alternative to separate optional supplemental instruction is incorporation of hints within the usual course assignments. Intelligent tutoring systems that students could use for homework have been in place for almost two decades now. These systems have been shown to be effective in improving student learning at the K-12 as well as college level for many disciplines such as mathematics [23, 24, 25, 26], accounting [27], and programming [28]. However, the development of these systems was time-intensive and required programming expertise. The emergence of Large language Models (LLMs) such as GPT-4 over the past couple of years has made this much easier [29]. For example, AI-generated hints are offered in all the online homework assignments on kudu.com. The hints are based on the text of the chapter the students are studying. The effectiveness of such hints generated by AI in supporting student learning are relatively unknown, but it holds much promise.
@@ I.1 Theoretical Framework: Expectancy-Value Theory of Achievement Motivation
In terms of expecting to do well, the AI-generated hints on course assignments could increase student expectations of success on the assignments. In terms of value, scaling extra credit points for the supplemental math assignments could increase the utility value of those assignments, especially for students that did not do well on the midterm exams. And given that the AI hints are integrated into the homework platform in Kudu and don’t take much additional time/effort, the cost is minimal.
AI-generated hints also reduce another aspect of the cost of getting help. Students often experience social barriers, including the fear of negative evaluation, that hinders them from asking instructors, teaching assistants, and/or peer for help [39, 40]. However, such social barriers do not exist when using AI tools. AI-generated hints might be particularly helpful for students from marginalized backgrounds that are severely underrepresented in the classroom and/or prone to stereotype threat, which is the term for “when members of a stigmatized group find themselves in a situation where negative stereotypes provide a possible framework for interpreting their behavior, the risk of being judged in light of those stereotypes can elicit a disruptive state that undermines performance and aspirations in that domain” [41].
@@ I.2 Research questions
  • RQ2: Does using optional supports increase student exam performance in an introductory Physics course?
  • RQ3: Is the impact of optional supports on student grades in the course similar for students across a range of social identities?
@@ I.3 Positionality of the authors
Y.L. is a graduate student in the Department of Physics and Astronomy at UCLA. He is working towards his PhD degree in physics and has been Teaching Assistants for multiple introductory and upper-level undergraduate physics courses. Y.L. identifies as an Asian male.
@@ II Methods
The study was carried out at a large public university in the Southwest United States. The students participating in the study were enrolled in one of two sections (hereafter referred to as section A and B) of a calculus-based introductory physics course on Mechanics, typically the first in a series of General Physics courses. Both sections were taught by the same instructor. We obtained demographic data from institutional records. All study protocols were approved by the IRB. The basic demographic composition of the participants is summarized in Fig. 1. To protect student privacy and identifiable information, we grouped Black/African American, Hispanic/Latine, and Native American students together for analyses because of small cell sizes. Following the recommendation from Asai, we call this grouping ”PEERs” i.e. persons excluded from STEM because of their ethnicity or race, because this term acknowledges that historical exclusion of people based on their race/ethnicity has been systematic and intentional [42].
@@ II.1 Study Design
We adopted a quasi-experimental design to assess the impact of optional supports on student exam performance. All supplemental materials and homework questions assigned to sections A and B were the same. The course was divided into three periods with one exam conducted at the end of each period. While math materials were available to students in both lectures, AI hints and extra credit for completing the math materials were only available to one of the two sections during each period (except in Period 3). The detailed intervention schedule is listed in Tab. 2.
To measure students’ prior preparation, we administered two concept inventories at the beginning of the course to measure students’ knowledge of Newtonian mechanics i.e. the Force Concept Inventory (FCI) [43] and Calculus Concept Inventory (CCI) [44].
@@ II.2 Optional supports
The AI-generated hints are qualitative explanations related to physics homework questions and often include key ingredients to solve problems. When a student requests a hint, the AI engine is provided with the text of the entire chapter and the text of the assignment, along with a carefully designed prompt instructing the AI to refrain from giving the full solution to the problem, while pointing the student in the right direction. An example of the AI-generated hints is demonstrated in App. B. We have carefully inspected the quality of the such hints and made sure that they are both relevant and beneficial to build up students’ problem solving skills in this course. When AI-generated hints are available, students have unlimited attempts to use them for each question so that they can explore the most useful version.
@@ II.3 Exam question evaluation
There were three exams used for summative assessment in the course, two midterms and a cumulative final. To disentangle the effect of AI-generated hints versus supplemental math material, we evaluated all exam questions to assess their alignment with the math included in the supplemental math assignments and Physics homework problems. The evaluation process was carried out as follows: 1) we selected a subset of homework questions that were relevant to the exam questions and created a question pool for each exam. 2) For each exam question, we determined if there was an isomorphic question (i.e. a question that is almost identical but with minor changes) in the question pool. 3) For each exam question, we also matched it to one or more supplemental math materials according to the math skills required to solve it. This evaluation was performed by Y.L. who has a bachelor’s degree in physics and has taught undergraduate physics courses.
@@ II.4 Regression models
We used multiple linear regression to assess the impact of optional supports and other variables on student exam performance. Student exam scores were used as the outcome variable. To standardize the data, we converted the exam scores into z-scores with zero mean and standard deviation 1, according to the following transformation
@@
where $\mu$ and $\sigma$ are the mean and standard deviation of each exam. The following factors were used as predictors: 1) prior preparation measured by the concept inventories and converted to z-scores, 2) homework scores, 3) optional support usage, including use of AI-generated hints and completion of supplemental math material, and 4) demographics, specifically gender and race/ethnicity. For each exam, we utilize the question pool from the exam evaluation stage (Sec. II.3) and quantify the AI-generated hints usage by the proportion of problems that a student used AI hints on among all the problems in the question pool. We note that each exam has a separate question pool, so that AI-generated hints usage is only measured on the relevant (but not necessarily isomorphic) homework questions. The race/ethnicity and gender variables were converted to binary variables to protect student privacy given some small cell sizes in the disaggregated data. White/Asian (WA) students were used as reference to PEER (persons excluded due to race/ethnicity including Black/African American, Hispanic/Latine, and Native American students, and men were used as reference to students with underrepresented gender identities (URG: women and non-binary). The details of all variables used in the model are listed in Tab. 2.
@@
where the dots include regression terms from other variables. When one of the variables in the interaction term ($x_{\mathrm{re}}$ here) is binary, the interaction coefficient has a simple interpretation: if the binary variable indicates the reference group (taking value 0), then the effect of the second variable is solely expressed in terms of its coefficient. However, when the binary variable takes the value 1, the effect of the second variable comes from the combination of its coefficient and the interaction coefficient. Therefore, in our example, $\beta_{int}$ indicates the difference between the effect of completing math supplemental materials between the race/ethnicity groups.
To quantify factors that affect students’ completion of math materials, we performed a logistic regression. Here, the outcome variable was whether or not students completed a given supplemental math assignment, and predictor variables included prior preparation, demographics, and which section students were enrolled in.
@@ III Results
In FCI, male students score an average of 0.24 (SD = 1.02) and WA students score an average of 0.10 (SD = 0.99), while URG students have mean = -0.30 (SD = 0.90) and PEERs have mean = -0.44 (SD = 0.91).
In CCI, male students score an average of 0.19 (SD = 0.93) and WA students score an average of 0.09 (SD = 0.98), while URG students have mean = -0.23 (SD = 1.04) and PEERs have mean = -0.37 (SD = 1.03).
III.1 Finding 1: Incentives increased completion and reduced disparities in completion of math supplemental material for PEERs
@@
On average, the completion rates for supplemental math assignments when unincentivized were around $20\%$ and offering extra credit increased the completion rate by about $20\%$. Furthermore, when incentives were not offered, we observed a significantly low completion rate for PEERs in Section B, compared to WA students (upper panel in Fig. 3). But such disparity was alleviated when extra credit was offered. Our full logistic regression results are summarized in App. C.
We observed no difference in the use of AI hints across gender and race/ethnicity groups Fig. 4. Most of the students explored the AI-generated hints feature in Kudu: when available, $88\%$ of the students used AI-generated hints on at least 1 homework question (among the problems that we found to match the exam questions), and $57\%$ students used AI-generated hints on at least 5 questions. Depending on the section, the questions pool size is around 40. On average, students used AI-generated hints on at least $20\%$ of the questions.
@@
Race/ethnicity: PEER -0.23 0.09 0.011
@@
Homework score 3.18 0.21 $<0.001$
AI hints usage 0.16 0.22 0.466
Math supplement material 1 0.30 0.08 $<0.001$
Math supplement material 2 -0.14 0.08 0.096
Math supplement material 3 0.08 0.08 0.293
Force concept inventory 0.31 0.04 $<0.001$
Calculus concept inventory 0.12 0.04 0.003
@@
The regression model for Exam 3 is shown in Tab. 3 and we list the results for Exam 1 and 2 in App. C. None of the exam problems used the math skills covered in supplemental math material 4, thus we excluded this variable from the model. We did not find a significant correlation between gender and exam performance. However, PEER students had significantly lower exam scores compared to WA students. The correlation between demographics and exam performance was consistent across all exams. In addition, prior preparation in physics/math had a significant association with student exam scores, which is consistent with the findings in Ref. [4].
@@
When we compared the relevance of material 1 indicated by our Relevance Index (RI, i.e. proportion of physics exam questions that used the math skills on a given math assignment) and its regression coefficient across all exams, we found that as RI increases, the regression coefficient becomes not only significant but also larger. This effect is visually demonstrated in Fig. 5, where we plot the best fit lines of exam score versus FCI score for students that completed supplemental math material 1 and those that did not. We observed a clear separation of the fitted lines for Exam 1 ($\rm RI=0.6$) and Exam 3 ($\rm RI=0.8$) with larger separation for Exam 3, but this is absent for Exam 2, which only has a RI of 0.2.
@@
We found a statistically significant association between Exam 2 score and AI hints usage, as demonstrated in Fig. 6 and Tab. 8. Interestingly, the linear fit result in Fig. 6 indicates that AI hints benefits students who are less prepared more significantly. To verify this effect, we added an interaction term between the AI hints usages and FCI variables, with the updated regression result shown in Tab. 9. Indeed, our observation is reflected in the negative coefficient of the interaction term. However, we did not find a significant association between student exam performance and AI hints usage for Exams 1 and 3, possibly due to the low relevance between these exams and the homework questions, as indicated from the ISI in Tab. 1.
III.4 Finding 3: Completing supplemental math assignments is associated with lower disparities in exam scores between WA students and PEERs
When we added an interaction between competing supplemental math material 1 and race/ethnicity in our regression model, we found that the interaction was statistically significant Tab. 4. The regression coefficients suggest that for two WAs with an average homework score and concept inventory score, the one who completed supplemental math assignment 1 would increase their exam score by about a quarter of a standard deviation. However, for two PEERs who had an average homework and concept inventory score, the one who completed the supplemental math assignment 1 would increase their exam score by more than half of a standard deviation, with contributions from both the math supplemental material 1 coefficient and the interaction term coefficient. Therefore, the additional math practice is not only beneficial to the entire class, but it also improves the exam performance of PEERs’ students even more than others.
@@
Homework score 3.16 0.21 $<0.001$
AI hints usage 0.17 0.22 0.457
Math supplement material 2 -0.13 0.08 0.122
Math supplement material 3 0.08 0.08 0.281
Force concept inventory 0.31 0.04 $<0.001$
Calculus concept inventory 0.13 0.04 0.002
Race/ethnicity -0.32 0.10 0.002
@@
Math supplement material 1 0.24 0.09 0.009
Race/ethnicity : Math material 1 0.36 0.20 0.066

### v2

# Incentivizing supplemental math assignments and using AI-generated hints is associated with improved exam performance

@@ Abstract
Inequities in student access to trigonometry and calculus are often associated with racial and socioeconomic privilege, and often influence introductory physics course performance. To mitigate these disparities in student preparedness, we developed a two-pronged intervention consisting of (1) incentivized supplemental math assignments and (2) AI-generated learning support tools in the form of optional hints embedded in the physics homework assignments. Both interventions are grounded in the Situated Expectancy-Value Theory of Achievement Motivation, which posits that students are more likely to complete a task that they expect to do well in and whose outcomes they think are valuable. For the supplemental math assignments, the extra credit was scaled to make it worth more points for students with lower exam scores, thereby creating even greater value for students who might benefit most from the assignments. AI-generated hints were integrated into the homework assignments, thereby reducing or eliminating the cost to the student, in terms of time, energy, and social barriers or fear of judgment. Our findings indicate that both these interventions are associated with increased exam scores; in particular, the scaled extra credit reduced disparities in completion of supplemental math assignments. These interventions, which are relatively simple for any instructor to implement, are therefore very promising for creating more equitable undergraduate quantitative courses.
@@ I Introduction
Students’ mathematical skills play an important role in shaping their learning and performance in undergraduate Introductory Physics courses [1, 2, 3, 4]. However, students enter these courses with wide variation in their prior exposure to advanced mathematics (i.e. trigonometry and calculus). Moreover, student access to advanced math courses is associated with racial and socioeconomic privilege, so that Black, Hispanic/Latine, and Native American students, and students from low-income backgrounds are less likely to have had the opportunity to take advanced math courses in high school [5, 6, 7, 8].
@@
One strategy to address the inequity in prior access to advanced mathematics courses is to provide supplemental instruction to students. This has been found to help students with less math preparation successfully complete Introductory Physics courses [15, 16, 4]. Most supplemental instruction takes the form of an optional in-person course or series of workshops focused on students’ problem solving skills in small groups [17, 16, 15]. This can be quite time-consuming and resource-intensive. Using online learning tools offers an alternative less resource-intensive way to approach supplemental instruction but relies more heavily on student motivation and might require students to pay additional fees [4, 18]. For example, Forrest et al. [4] offered an optional asynchronous online math tutorial to Introductory Physics students who scored less than 65% on a Math diagnostic exam and found that students who completed the online tutorial were four times more likely to pass the course. However, this tutorial was not free and only about half of the students completed the online tutorial, even though completing the tutorial was associated with a small amount of course points [4]. Thus, motivating students to participate in optional online supplemental math instruction can be challenging [19, 18].
@@
An alternative to separate optional supplemental instruction is incorporation of hints within the usual course assignments. Intelligent tutoring systems that students could use for homework have been in place for almost two decades now. These systems have been shown to be effective in improving student learning at the K-12 as well as college level for many disciplines such as mathematics [23, 24, 25, 26], accounting [27], and programming [28]. However, the development of these systems was time-intensive and required programming expertise. The emergence of Large-language Models (LLMs) such as GPT-4 over the past couple of years has made this much easier [29]. For example, AI-generated hints are offered in all the online homework assignments on kudu.com. The hints are based on the text of the chapter the students are studying. The effectiveness of such hints generated by AI in supporting student learning are relatively unknown, but it holds much promise.
@@ I.1 Theoretical Framework: Expectancy-Value Theory of Achievement Motivation
In terms of expecting to do well, the AI-generated hints on course assignments could increase student expectations of success on the assignments. In terms of value, scaling extra credit points for the supplemental math assignments could increase the utility value of those assignments, especially for students that did not do well on the midterm exams. And given that the AI-generated hints are integrated into the homework platform in Kudu and don’t take much additional time/effort, the cost is minimal.
AI-generated hints also reduce another aspect of the cost of getting help. Students often experience social barriers, including the fear of negative evaluation, that hinders them from asking instructors, teaching assistants, and/or peers for help [39, 40]. However, such social barriers do not exist when using AI tools. AI-generated hints might be particularly helpful for students from marginalized backgrounds that are severely underrepresented in the classroom and/or prone to stereotype threat, which is the term for “when members of a stigmatized group find themselves in a situation where negative stereotypes provide a possible framework for interpreting their behavior, the risk of being judged in light of those stereotypes can elicit a disruptive state that undermines performance and aspirations in that domain” [41].
@@ I.2 Research questions
  • RQ2: Is using optional supports associated with an increase in student exam scores?
  • RQ3: Is the use of optional supports associated with student grades similarly for students across a range of social identities?
@@ I.3 Positionality of the authors
Y.L. is a graduate student in the Department of Physics and Astronomy at UCLA. He is working towards his PhD degree in physics and has been Teaching Assistant for multiple introductory and upper-level undergraduate physics courses. Y.L. identifies as an Asian male.
@@ II Methods
The study was carried out at a large public university in the Southwest United States. The students participating in the study were enrolled in one of two sections (hereafter referred to as section A and B) of a calculus-based introductory physics course on Mechanics, typically the first in a series of General Physics courses. Both sections were taught by the same instructor. We obtained demographic data from institutional records. All study protocols were approved by the UCLA IRB. The basic demographic composition of the participants is summarized in Fig. 1. To protect student privacy and identifiable information, we grouped Black/African American, Hispanic/Latine, and Native American students together for analyses because of small cell sizes. Following the recommendation from Asai, we call this grouping ”PEERs” i.e. persons excluded from STEM because of their ethnicity or race, because this term acknowledges that historical exclusion of people based on their race/ethnicity has been systematic and intentional [42].
@@ II.1 Study Design
We adopted a quasi-experimental design to assess the association between use of optional supports and student exam performance. All supplemental materials and homework questions assigned to sections A and B were the same. The course was divided into three periods with one exam conducted at the end of each period. While math materials were available to students in both lectures, AI-generated hints and extra credit for completing the math materials were only available to one of the two sections during each period (except in Period 3). The detailed intervention schedule is listed in Fig. 2.
To measure students’ prior preparation, we administered two concept inventories at the beginning of the course to measure students’ knowledge of Newtonian mechanics i.e. the Force Concept Inventory (FCI) [43] and Calculus Concept Inventory (CCI) [44]. Throughout the three periods of this study, we collected data on student homework scores, use of AI-generated hints on the homework problems, and completion of supplemental math materials. We combined this data with demographics to address RQ1 (see Sec. III.1). We then performed multiple linear regression with student exam performance as the outcome variable. This allowed us to quantitatively investigate RQ2 and RQ3 (see Sec. III.2, Sec. III.2 and Sec. III.3).
@@ II.2 Optional supports
The AI-generated hints are qualitative explanations related to physics homework questions and often include key ingredients to solve problems. When a student requests a hint, the AI engine is provided with the text of the entire chapter and the text of the assignment, along with a carefully designed prompt instructing the AI to refrain from giving the full solution to the problem, while pointing the student in the right direction. An example of the AI-generated hints is demonstrated in App. B. We have carefully inspected the quality of such hints and made sure that they are both relevant and beneficial to build up students’ problem solving skills in this course. When AI-generated hints are available, students have unlimited attempts to use them for each question so that they can explore the most useful version.
@@ II.3 Exam question evaluation
There were three exams used for summative assessment in the course, two midterms and a cumulative final. To disentangle the effect of AI-generated hints versus supplemental math material, we evaluated all exam questions to assess their alignment with the math included in the supplemental math assignments and Physics homework problems. The evaluation process was carried out as follows: 1) we selected a subset of homework questions that were relevant to the exam questions and created a question pool for each exam. 2) For each exam question, we determined if there was an isomorphic question (i.e. a pair of questions that require the same set of concepts and skills to solve, with differences only at the surface level, such as the provided numerical values or the problem context) in the question pool. 3) For each exam question, we also matched it to one or more supplemental math materials according to the math skills required to solve it. An example of a pair of isomorphic questions is given in App. C. This evaluation was performed by Y.L. who has a bachelor’s degree in physics and has taught undergraduate physics courses.
@@ II.4 Regression models
We used multiple linear regression to assess the association between optional supports and student exam performance. Student exam scores were used as the outcome variable. To standardize the data, we converted the exam scores into z-scores with zero mean and standard deviation 1, according to the following transformation
@@
where $\mu$ and $\sigma$ are the mean and standard deviation of each exam. The following variables were used as predictors: 1) prior preparation measured by the concept inventories and converted to z-scores, 2) homework scores, 3) optional support usage, including use of AI-generated hints and completion of supplemental math material, and 4) demographics, specifically gender and race/ethnicity. For each exam, we used the question pool from the exam evaluation stage (Sec. II.3) and quantified AI-generated hints use as the proportion of problems that a student used AI-generated hints on among all the problems in the question pool. We note that each exam has a separate question pool, so that AI-generated hints use is only measured on the relevant (but not necessarily isomorphic) homework questions. The race/ethnicity and gender variables were converted to binary variables to protect student privacy given some small cell sizes in the disaggregated data. White/Asian (WA) students were used as reference to PEER (persons excluded due to race/ethnicity including Black/African American, Hispanic/Latine, and Native American students, and men were used as reference to students with underrepresented gender identities (URG: women and non-binary). The details of all variables used in the model are listed in Tab. 2.
@@
where the dots indicate regression terms from other variables. When one of the variables in the interaction term ($x_{\mathrm{re}}$ here) is binary, the interaction coefficient has a simple interpretation: if the binary variable indicates the reference group (taking value 0), then the effect of the second variable is solely expressed in terms of its coefficient. However, when the binary variable takes the value 1, the effect of the second variable comes from the combination of its coefficient and the interaction coefficient. Therefore, in our example, $\beta_{int}$ indicates the difference between the effect of completing math supplemental materials between the race/ethnicity groups.
When we tested the assumptions for multiple linear regressions, we found that there is heteroscedasticity in our data. Therefore, to reliably estimate the $p$ values, we adopted the heteroscedasticity-consistent covariance ’HC3’ [45] when performing the regression. In addition, some variables in Tab. 2 are strongly correlated. As an example, Math supplement material 1 and 2 are positively correlated as they are incentivized together. We remedied the multicollinearity issue by dropping some of the variables when necessary (see Sec. III.2 for more details).
To quantify factors that are associated with students’ completion of math materials, we performed a logistic regression. Here, the outcome variable was whether or not students completed a given supplemental math assignment, and predictor variables included prior preparation, demographics, and which section students were enrolled in.
@@ III Results
Male students scored an average of 0.24 (SD = 1.02) on the FCI compared to URG students who had an average of -0.30 (SD = 0.90). WA students scored an average of 0.10 (SD = 0.99), while PEER students had mean of -0.44 (SD = 0.91).
On the CCI, male students scored an average of 0.19 (SD = 0.93) while URG students had mean of -0.23 (SD = 1.04). WA students scored an average of 0.09 (SD = 0.98) compared to PEER students who had mean of -0.37 (SD = 1.03).
III.1 Finding 1: Incentives are associated with higher completion rate and reduced disparities in completion of math supplemental material for PEERs
@@
On average, the completion rates for supplemental math assignments when unincentivized were around $20\%$ and offering extra credit increased the completion rate by about $20\%$. Furthermore, when incentives were not offered, we observed a significantly low completion rate for PEERs in Section B, compared to WA students (upper panel in Fig. 3). But this disparity was alleviated when extra credit was offered. Our full logistic regression results are summarized in App. D.
We observed no difference in the use of AI-generated hints across gender and race/ethnicity groups Fig. 4. Most of the students explored the AI-generated hints feature in Kudu: when available, $88\%$ of the students used AI-generated hints on at least 1 homework question (among the problems that we found to match the exam questions), and $57\%$ students used AI-generated hints on at least 5 questions. Depending on the section, the question pool size is around 40. On average, students used AI-generated hints on at least $20\%$ of the questions.
@@
Race/ethnicity: PEER -0.22 0.10 0.030
@@
Homework score 3.16 0.35 $<0.001$
AI hints usage 0.24 0.21 0.240
Math supplement material 1 0.24 0.07 0.001
Math supplement material 3 0.05 0.07 0.444
Force concept inventory 0.31 0.06 $<0.001$
Calculus concept inventory 0.12 0.06 0.029
@@
The regression model for Exam 3 is shown in Tab. 3 and we list the results for Exam 1 and 2 in App. D. None of the exam problems used the math skills covered in supplemental math material 4, thus we excluded this variable from the model. Additionally, there was a strong correlation between math material 1 and 2 since they were incentivized together, therefore we dropped the math material 2 variable from our models. We did not find a significant correlation between gender and exam performance. However, PEER students had significantly lower exam scores compared to WA students. The correlation between demographics and exam performance was consistent across all exams. In addition, prior preparation in physics/math had a significant association with student exam scores, which is consistent with the findings in Ref. [4].
@@
When we compared the relevance of material 1 indicated by our Relevance Index (RI, i.e. proportion of physics exam questions that used the math skills on a given math assignment) and its regression coefficient across all exams, we found that as RI increases, the regression coefficient becomes not only significant but also larger (see Tab. 3, Tab. 7, and Tab. 8). This effect is visually demonstrated in Fig. 5, where we plotted the best fit lines of exam score versus FCI score for students that completed supplemental math material 1 and those that did not. We observed a clear separation of the fitted lines for Exam 1 ($\rm RI=0.6$) and Exam 3 ($\rm RI=0.8$) with larger separation for Exam 3, but this is absent for Exam 2, which only has a RI of 0.2.
@@
We found a statistically significant association between Exam 2 score and AI hints usage, as demonstrated in Fig. 6 and Tab. 8. Interestingly, the linear fit result in Fig. 6 suggests that AI-generated hints might benefit students who are less prepared to a larger extent. To examine this effect, we added an interaction term between the AI hints usages and FCI variables, with the updated regression result shown in Tab. 9. Indeed, our observation is reflected in the negative coefficient of the interaction term, although not statistically significant. However, we did not find a significant association between student exam performance and AI hints usage for Exams 1 and 3, possibly due to the low relevance between these exams and the homework questions, as indicated from the ISI in Tab. 1.
III.4 Finding 3: Completing supplemental math assignments might be associated with lower disparities in exam scores between WA students and PEERs
When we added an interaction between competing supplemental math material 1 and race/ethnicity in our regression model, we found that the interaction coefficient was positive with a large effect size (coefficient of 0.38, compared to a coefficient of 0.32 for the Force Concept Inventory) and a $p$ value of 0.081 (see Tab. 4). Although the $p$ value does not reach the conventional threshold at 0.05, the interaction term still suggests that completing supplemental math materials might benefit PEER students more.
The regression coefficients suggest that for two WA students with an average homework score and concept inventory score, the one who completed supplemental math assignment 1 would increase their exam score by about a fifth of a standard deviation. However, for two PEER students who had an average homework and concept inventory score, the one who completed the supplemental math assignment 1 would increase their exam score by more than half of a standard deviation, with contributions from both the math supplemental material 1 coefficient and the interaction term coefficient. Therefore, our result provides some tentative evidence that the additional math practice is not only beneficial to the entire class, but it also improves the exam performance of PEERs’ students even more than others.
@@
Homework score 3.14 0.35 $<0.001$
AI hints usage 0.24 0.21 0.245
Math supplement material 3 0.06 0.07 0.414
Force concept inventory 0.32 0.06 $<0.001$
Calculus concept inventory 0.12 0.06 0.023
Race/ethnicity -0.32 0.12 0.006
@@
Math supplement material 1 0.18 0.08 0.020
Race/ethnicity : Math material 1 0.38 0.22 0.081